Morrison Company uses a job-order costing system to assign manufacturing costs to jobs. Its balance sheet on January 1 is as follows:

Morrison Company
Balance Sheet
January 1
Assets
Cash $ 45,900
Raw materials $ 16,900
Work in process 5,450
Finished goods 18,600 40,950
Prepaid expenses 3,500
Property, plant, and equipment (net) 145,000
Total assets $ 235,350
Liabilities and Stockholders’ Equity
Accounts payable $ 7,800
Retained earnings 227,550
Total liabilities and stockholders’ equity $ 235,350

During January the company completed the following transactions:

  1. Purchased raw materials on account, $78,200.
  2. Raw materials used in production, $91,600 ($77,000 was direct materials and $14,600 was indirect materials).
  3. Paid $176,750 of salaries and wages in cash ($95,600 was direct labor, $37,500 was indirect labor, and $43,650 was related to employees responsible for selling and administration).
  4. Various manufacturing overhead costs incurred (on account) to support production, $39,900.
  5. Depreciation recorded on property, plant, and equipment, $72,000 (70% related to manufacturing equipment and 30% related to assets that support selling and administration).
  6. Various selling expenses paid in cash, $33,650.
  7. Prepaid insurance expired during the month, $2,200 (80% related to production, and 20% related to selling and administration).
  8. Manufacturing overhead applied to production, $134,000.
  9. Cost of goods manufactured, $295,200.
  10. Cash sales to customers, $402,280.
  11. Cost of goods sold (unadjusted), $290,200.
  12. Cash payments to creditors, $71,200.
  13. Underapplied or overapplied overhead  $?  .

Calculate the ending balances that would be reported on the company's balance sheet on January 31st. (Hint: Be sure to calculate the underapplied or overapplied overhead and then account for its affect on the balance sheet.) (Amounts to be deducted should be indicated by a minus sign.)

Morrison Company
Transaction Analysis
For the Month Ended Jaunary 31
Transactions Cash Raw Materials Work in Process Finished Goods Manufacturing Overhead Prepaid Expenses PP&E (net) = Accounts Payable Retained Earnings
Beginning balances @1/1 $45,900 $16,900 $5,450 $18,600 $0 $3,500 $145,000 = $7,800 $227,550
(a) Raw material purchases 78,200 =
(b) Raw materials used in production =
(c) Salaries and wages =
(d) Various overhead costs =
(e) Depreciation =
(f) Various selling expenses =
(g) Expiration of prepaid insurance =
(h) Manufacturing overhead applied =
(i) Cost of goods manufactured =
(j) Sales =
(k) Cost of goods sold =
(l) Payments to creditors =
(m) =
Ending balances @ 1/31 $73,000
  • Required 1

Required:

1. Calculate the ending balances that would be reported on the company's balance sheet on January 31st. (Hint: Be sure to calculate the underapplied or overapplied overhead and then account for its affect on the balance sheet.)

2. What is Morrison Company’s net operating income for the month of January?
Answer #1

1.

Morrison Company
Transaction Analysis
For the Month Ended Jaunary 31
Transactions Cash Raw Materials Work in Process Finished Goods Manufacturing Overhead Prepaid Expenses PP&E (net) = Accounts Payable Retained Earnings
Beginning balances @1/1 $45,900 $16,900 $5,450 $18,600 $0 $3,500 $145,000 = $7,800 $227,550
(a) Raw material purchases 78,200 = 78,200
(b) Raw materials used in production -91,600 77,000 14,600 =
(c) Salaries and wages -176,750 95,600 37,500 = -43,650
(d) Various overhead costs 39,900 = 39,900
(e) Depreciation 50,400 -72,000 = -21,600
(f) Various selling expenses -33,650 = -33,650
(g) Expiration of prepaid insurance 1,760 -2,200 = -440
(h) Manufacturing overhead applied 134,000 -134,000 =
(i) Cost of goods manufactured -295,200 295,200 =
(j) Sales 402,280 = 402,280
(k) Cost of goods sold -290,200 = -290,200
(l) Payments to creditors -71,200 = -71,200
(m) Underapplied overhead -10,160 = -10,160
Ending balances @ 1/31 $166,580 $3,500 $16,850 $23,600 $0 $1,300 $73,000 $54,700 $230,130

______________________________________________________________________________________

2.

Morrison Company’s net operating income for the month of January

Particulars Amount
Sales $402,280
Less: COGS ($290,200)
Less: Underapplied overhead ($10,160)
Gross profit $101,920
Less: Expenses
Salaries & wages expense ($43,650)
Depreciation expense ($21,600)
Selling expense ($33,650)
Insurance expense ($440) ($99,340)
Net Income $2,580